OpenStudy (anonymous):

x-8 =2 solve the equation for x,

OpenStudy (angela210793):

add 8 to both sides

OpenStudy (anonymous):

x-8+8=2+8 x= 10 ... Havent you got the hang of this by now lol (:

OpenStudy (anonymous):

x - 8 =2 +8 +8 (add 8 to both sides) x = 10

OpenStudy (anonymous):

let's say you have x number of lollipops and you need to take away 8 from it so that it equals 2 all you have to do now is figure out what equals 8 +2 and then you have x